# Simulation Based Bayesian Optimization

### v1

@@ Abstract
Bayesian Optimization (BO) is a powerful method for optimizing black-box functions by combining prior knowledge with ongoing function evaluations. BO constructs a probabilistic surrogate model of the objective function given the covariates, which is in turn used to inform the selection of future evaluation points through an acquisition function. For smooth continuous search spaces, Gaussian Processes (GPs) are commonly used as the surrogate model as they offer analytical access to posterior predictive distributions, thus facilitating the computation and optimization of acquisition functions. However, in complex scenarios involving optimizations over categorical or mixed covariate spaces, GPs may not be ideal.
This paper introduces Simulation Based Bayesian Optimization (SBBO) as a novel approach to optimizing acquisition functions that only requires sampling-based access to posterior predictive distributions. SBBO allows the use of surrogate probabilistic models tailored for combinatorial spaces with discrete variables. Any Bayesian model in which posterior inference is carried out through Markov chain Monte Carlo can be selected as the surrogate model in SBBO. In applications involving combinatorial optimization, we demonstrate empirically the effectiveness of SBBO method using various choices of surrogate models.
Keywords: Acquisition functions, Bayesian decision theory, Black-box functions, Combinatorial optimization, Markov chain Monte Carlo
@@ 1 Introduction
The problem of optimizing a potentially nonlinear function $f(\boldsymbol{x})$ over a certain set $\mathcal{S}$ has been extensively studied. In many applications, the objective function is not explicitly known. This can make optimization challenging, as it is difficult to determine the optimal solution without accurate knowledge of the function. Moreover, $f(\boldsymbol{x})$ may be expensive to evaluate and a budget constraint often restricts evaluations of the objective function to a sparse selection of points, further complicating the optimization task. As an example, consider the process of drug discovery. During its early stages, the goal is to identify candidate molecules with a desired property, such as maximum binding affinity to a given biological target. However, measuring this binding affinity requires synthesizing and testing molecules in vitro, limiting the number of leads that can be evaluated. Furthermore, expressing affinity as a mathematical function of the molecule’s structure is not possible.
@@
Bayesian Optimization (BO, Mockus (1994)) stands as an efficient approach for identifying the extrema of a black-box objective function. Its effectiveness originates from its capacity to integrate prior knowledge about the function with the evaluations previously conducted, enabling informed decisions on where to carry out additional evaluations. Bayesian Optimization involves the construction of a Bayesian surrogate model for $f(\boldsymbol{x})$. Subsequently, the next evaluation point is selected through the maximization of an acquisition function (Frazier, 2018). Commonly used acquisition functions, including expected improvement, are defined as the expected values of utility functions computed with respect to the posterior predictive distribution of $f(\boldsymbol{x})$ derived from the Bayesian surrogate model. In the remainder of the paper, we will restrict our attention to this class of acquisition functions.
@@
Much of the BO literature focuses on continuous search spaces, where Gaussian Processes (GPs, Rasmussen et al. (2006)) are predominantly chosen as the Bayesian surrogate model. GPs are well-suited to optimization tasks in continuous search spaces as they rely on the smoothness defined by a kernel to model uncertainty. Moreover, GPs offer the advantage of yielding a analytical posterior predictive distribution (the posterior is Gaussian), enabling the analytical computation of commonly used acquisition functions like expected improvement or probability of improvement. However, in more intricate scenarios, such as those involving combinatorial (i.e. categorical) or mixed categorical-continuous covariate spaces (i.e. where the search space is a combination of discrete and continuous variables), GPs may not be the most suitable choice.
Related work.
In the last few years there has been some research in how to perform Bayesian optimization over combinatorial domains. As highlighted by Wang et al. (2023), this task comes with two major challenges: the construction of effective surrogate models over the combinatorial space, and the effective search in the combinatorial domain for the next structure for evaluation according to the acquisition function. Most previous research has been devoted to the first task. Within this, an important line of investigation consists of modifying GP kernel functions to accommodate categorical covariate spaces. For instance Garrido-Merchán and Hernández-Lobato (2020) treat categorical variables as continuous via one-hot encoding. Then, within the kernel function, continuous covariates are suitably transformed to one-hot encodings. Oh et al. (2019) constructs a combinatorial graph over the combinatorial search space. Subsequently, graph Fourier transforms are used to derive a diffusion kernel on the graph.
@@
Another interesting line of research performs BO over a continuous latent representation of the combinatorial spaces learned through a Variational Autoencoder (Gómez-Bombarelli et al., 2018). Finally, a third line of research gets rid of the GPs, and proposes using inherently discrete surrogate models. For instance Hutter et al. (2011) use random forests to model the objective function given the covariates. Baptista and Poloczek (2018) model the response as a linear function of the covariates including all possible pairwise interactions. A heavy-tailed horseshoe prior is used to induce sparsity on the coefficients.
In this paper, we focus on the second challenge: the task of effective search within the combinatorial domain. We introduce Simulation Based Bayesian Optimization (SBBO), a novel approach to optimize acquisition functions in Bayesian Optimization. Unlike previous approaches, SBBO does not require knowing the posterior predictive distribution analytically. All that is needed is sampling-based access to posterior predictive distributions. This allows us to use any Bayesian model whose posterior inference is done via Markov chain Monte Carlo (MCMC). Thus, determining where to make the next function evaluation becomes a simpler task that does not rely on closed-form results that are typically only available under the Gaussian surrogate. SBBO opens the door to using surrogate probabilistic models that may be better tailored for Bayesian optimization tasks in combinatorial spaces, where the search for optimal solutions becomes more challenging due to the presence of discrete variables.
@@
The paper is structured as follows: Section 2 provides an overview of Bayesian Optimization. Section 3 presents SBBO and summarises the surrogate probabilistic models that we use for the experiments presented in Section 4. Finally, Section 5 gives some concluding remarks and lines of future research.
@@ 2 Bayesian Optimization
In the setting of Bayesian optimization, the objective function $f(\boldsymbol{x})$ is assumed to be a black box, so no assumptions of convexity or linearity may be made. Additionally, it is assumed that $f$ is expensive to evaluate in that the number of evaluations of $f$ that can be feasibly performed is small, perhaps due to computing time or financial cost. It is further assumed that querying the function at $x$ provides a noisy observation $y$ of $f(x)$, with normal zero mean, $\sigma^{2}$ variance noise.
BO uses a surrogate probabilistic model of the objective function, usually obtained through Bayesian methods, to inform the next evaluation decision. It works in an iterative algorithm by 1) fitting the surrogate model and 2) using an acquisition function to obtain a new point $\boldsymbol{x}$ at which we evaluate or observe $f(\boldsymbol{x})$. Specifically, the surrogate model is fitted using all available data $\mathcal{D}_{1:n}=\{(\boldsymbol{x}_{i},y_{i}),i=1,\ldots,n\}$. Then using an acquisition function, we select a new $\boldsymbol{x}_{n+1}$ and obtain a corresponding $y_{n+1}$. The surrogate model is then updated to include the additional data point $(\boldsymbol{x}_{n+1},y_{n+1})$. We briefly review each of these two key components of Bayesian Optimization: the surrogate model and the acquisition function.
The following observational model is assumed
@@
where $f(\boldsymbol{x}_{t})$ could be parametric or non-parametric. Given a new location $\boldsymbol{x}$ and appropriate choice of prior, the posterior predictive distribution on $f(\boldsymbol{x}_{t})$ is taken as the surrogate model. If a parametric model is chosen for $f$, then $f(\boldsymbol{x}_{t})$ will depend on parameters $\boldsymbol{\beta}$ and is denoted as $f_{\boldsymbol{\beta}}(\boldsymbol{x}_{t})$. A prior $\pi(\boldsymbol{\beta})$ for the parameters $\boldsymbol{\beta}$ needs to be specified. Coupled with a likelihood $\pi(y|\boldsymbol{x},\boldsymbol{\beta},\sigma^{2})$, after observing data $\mathcal{D}_{1:n}$, we obtain the posterior $\pi(\boldsymbol{\beta}|\mathcal{D}_{1:n})$. Then, the surrogate model is
$\pi(f(\boldsymbol{x})|\boldsymbol{x},\mathcal{D}_{1:n})\equiv\pi(f_{\boldsymbol{\beta}}(\boldsymbol{x})|\boldsymbol{x},\mathcal{D}_{1:n})=\mathbb{E}_{\boldsymbol{\beta}|\mathcal{D}_{1:n}}\left[f_{\boldsymbol{\beta}}(\boldsymbol{x})\right]$
@@
An acquisition function is used to direct the search for the optimal $\boldsymbol{x}^{*}\in\mathcal{S}$. There are several possible choices. In this paper we restrict our attention to the class of acquisition functions that can be written as expectations of a utility function with respect to the posterior $\pi(f(\boldsymbol{x})|\boldsymbol{x},\mathcal{D}_{1:n})$. Some of the most commonly used acquisition functions belong to this class, such as expected improvement and probability of improvement. Let $u(f(\boldsymbol{x}),\boldsymbol{x})$ denote the utility achieved by evaluating the function at a new point $\boldsymbol{x}$ and obtaining $f(\boldsymbol{x})$.
Thus, the next point $\boldsymbol{x}_{n+1}$ is chosen as the one that maximizes the acquisition function, i.e. the expectation of the utility function:
$\boldsymbol{x}_{n+1}=\arg\max_{\boldsymbol{x}\in\mathcal{S}}\mathbb{E}_{f|\boldsymbol{x},\mathcal{D}_{1:n}}\left[u(f(\boldsymbol{x}),\boldsymbol{x})\right].$ (2.3)
Under the choice of GP prior for $f$, the posterior is Gaussian and under common choices of the utility functions, there is a closed-form expression for the expectation (2.3). The BO framework is indeed flexible, but computationally efficient and analytic results rely on this choice of Gaussian process prior. However, in specific scenarios-such as when the search space comprises categorical or mixed categorical-continuous variables—this GP prior may not be the best fit. In these situations, various alternative Bayesian models better tailored to the problem may be available. However, the alternatives often lack an analytically-defined posterior distribution, obstructing the selection of the next point $\boldsymbol{x}_{n+1}$ according to (2.3). We note, however, that we can sample from the posterior. To address this challenge presented by a non-GP choice of prior, we introduce a Bayesian optimization methodology that exclusively requires sampling access to the posterior, enabling the use of Bayesian models better suited to structured covariates spaces.
It is important to emphasize that choosing the next sampling location in (2.3) implicitly entails adopting a “one-step” assumption: the selection is made under the presumption that no future evaluations will take place. When there is budget for more than one function evaluation, Expected Improvement optimization unequivocally adopts a myopic approach. While several alternative strategies have been proposed for multi-step lookahead Bayesian Optimization (Wu and Frazier, 2019; Lam et al., 2016), we focus on the one-step look-ahead case.
@@ 3 Simulation Based Bayesian Optimization
We incorporate ideas from the simulation-based optimal Bayesian design literature (Müller, 2005) into the field of Bayesian Optimization, yielding our proposed method of Simulation-Based Bayesian Optimization (SBBO).
@@ 3.1 Main Idea
As in all Bayesian Optimization tasks, we aim to solve the optimization problem 2.1. The goal is to estimate the objective function $f$ with few evaluations. We do not assume that the objective function $f$ is continuous, but otherwise all the difficulties of estimating $f$ hold.
Imagine we have access to (perhaps through previous experiments or function evaluations) data $\mathcal{D}_{1:n}=\{(\boldsymbol{x}_{i},y_{i}),i=1,\ldots,n\}$, where $\boldsymbol{x}_{i}$ is a vector of covariates for observation $i$ and $y_{i}$ is obtained from the observation model in (2.2). Suppose that we have a budget of at most $M$ more function evaluations. Using all of the available information, the task is to determine the next $\boldsymbol{x}_{n+1},\boldsymbol{x}_{n+2},\ldots,\boldsymbol{x}_{n+M}$ at which we will evaluate the objective function. As is common practice in BO, rather than deciding all $M$ new locations/vectors at once, we proceed in a greedy manner by deciding on new locations in an iterative method.
To select the next point, $\boldsymbol{x}_{n+1}$, we typically maximize an acquisition function computed using $\pi(f(\boldsymbol{x})|\boldsymbol{x},\mathcal{D}_{1:n})$ over possible values of $\boldsymbol{x}$. Let $u(f(\boldsymbol{x}),\boldsymbol{x})$ be a non-negative and bounded utility function that quantifies the benefits/utility/gains from evaluating the objective function at $\boldsymbol{x}$ and observing the corresponding $f(\boldsymbol{x})$, assuming we have already observed data $\mathcal{D}_{1:n}$. Following Bayesian decision-theoretic principles, the next evaluation should be performed at the location $\boldsymbol{x}$ that maximizes the expected utility
[…]
@@
In order to perform Bayesian optimization, all we need is to (1) decide on a suitable Bayesian surrogate model and (2) solve (3.1). As previously mentioned, except in special cases, the posterior $\pi(f(\boldsymbol{x})|\boldsymbol{x},\mathcal{D}_{1:n})$ is typically only accessible through sampling. For this purpose, we borrow ideas from the optimal Bayesian design community. The key idea is to reframe optimization of (3.1) as a simulation problem. Following Bielza et al. (1999), we define an augmented distribution $g$ that incorporates $\boldsymbol{x}$ as a random variable:
$g(\boldsymbol{x},f(\boldsymbol{x})|\mathcal{D}_{1:n})\propto u(f(\boldsymbol{x}),\boldsymbol{x})\pi(f(\boldsymbol{x})|\boldsymbol{x},\mathcal{D}_{1:n}).$ (3.2)
@@
It is straightforward to see that the marginal of $g$ in $\boldsymbol{x}$ is proportional to the expected utility defined in (3.1). Thus, the mode of this marginal distribution coincides with the optimal next evaluation point. This suggests a simulation-based approach to find $\boldsymbol{x}_{n+1}$: we can simply simulate from $(\boldsymbol{x},f(\boldsymbol{x}))\sim g(\boldsymbol{x},f(\boldsymbol{x})|\mathcal{D}_{1:n})$ using standard MCMC techniques and compute the mode of $g(\boldsymbol{x}|\mathcal{D}_{1:n})$ using the generated samples $\boldsymbol{x}$. However, this strategy is limited to low-dimensional $\boldsymbol{x}$. For high-dimensional covariate spaces, estimating the mode of $g(\boldsymbol{x}|\mathcal{D}_{1:n})$ from the simulation output becomes impracticable. Müller et al. (2004) propose an alternative augmented distribution $g_{H}$ defined as
@@
where $f(\boldsymbol{x})_{1},\ldots,f(\boldsymbol{x})_{H})$ are $H$ copies of the random variable $f(\boldsymbol{x})$. Note that the marginal of $g_{H}$ in $\boldsymbol{x}$ is
$g_{H}(\boldsymbol{x}|\mathcal{D}_{1:n})\propto\Psi(\boldsymbol{x})^{H}.$ (3.4)
Thus, the mode of this marginal distribution also coincides with the optimal point at which we should make our next function evaluation. In particular, if $H$ is sufficiently large, the marginal distribution $g_{H}(\boldsymbol{x})$ is tightly concentrated on the next $\boldsymbol{x}_{n+1}$.
Müller et al. (2004) propose an inhomogeneous MCMC sampling scheme to simulate from $g_{H}$. They begin by generating samples from $g$ using the standard Metropolis-Hastings algorithm. In subsequent iterations, the target distribution is changed to $g_{H}$, with $H$ increasing across iterations. The inhomogeneous MCMC is constructed in such a way that the stationary distribution for fixed $H$ is $g_{H}$. We use this idea to find the next evaluation point in the context of Bayesian optimization. The algorithm is illustrated in Algorithm 1, where a Metropolis-Hastings sampling scheme is used.
@@
Obtain initial dataset $\mathcal{D}_{1:n}=\{(\boldsymbol{x}_{i},y_{i}),i=1,\ldots,n\}$. Determine form of utility function $u_{\mathcal{D}}(\boldsymbol{x},y)$.
@@
    Choose initial $\boldsymbol{x}_{n+1}^{(0)}$. For $h=1,\ldots,H$, simulate $f(\boldsymbol{x})^{(0)}_{h}\sim\pi(f(\boldsymbol{x})|\boldsymbol{x}_{n+1}^{(0)},\mathbf{D}_{1:n})$ and evaluate corresponding utility $u^{(0)}_{h}=u(f(\boldsymbol{x}_{n+1}^{(0)}),\boldsymbol{x}_{n+1}^{(0)})$. Define $v^{(0)}=\frac{1}{H}\sum_{h=1}^{H}\log u^{(0)}_{h}$.
@@
    For $h=1,\ldots,H$, simulate $\tilde{f}(\boldsymbol{x})_{h}\sim\pi(f(\boldsymbol{x})|\tilde{\boldsymbol{x}}_{n+1},\mathcal{D}_{1:n})$ and evaluate corresponding utility $\tilde{u}_{h}=u(\tilde{f}(\boldsymbol{x})_{h},\tilde{\boldsymbol{x}}_{n+1})$. Define $\tilde{v}=\frac{1}{H}\sum_{h=1}^{H}\log\tilde{u}_{h}$.
@@
    $\alpha$ $=\min\left\{1,\exp(H\tilde{v}-Hv^{(0)})\frac{q(\boldsymbol{x}^{(0)}_{n+1}|\tilde{\boldsymbol{x}}_{n+1})}{q(\tilde{\boldsymbol{x}}_{n+1}|\boldsymbol{x}^{(0)}_{n+1})}\right\}$

### v2

@@ Abstract
Bayesian Optimization (BO) is a powerful method for optimizing black-box functions by combining prior knowledge with ongoing function evaluations. BO constructs a probabilistic surrogate model of the objective function given the covariates, which is in turn used to inform the selection of future evaluation points through an acquisition function. For smooth continuous search spaces, Gaussian Processes (GPs) are commonly used as the surrogate model as they offer analytical access to posterior predictive distributions, thus facilitating the computation and optimization of acquisition functions. However, in complex scenarios involving optimization over categorical or mixed covariate spaces, GPs may not be ideal. This paper introduces Simulation Based Bayesian Optimization (SBBO) as a novel approach to optimizing acquisition functions that only requires sampling-based access to posterior predictive distributions. SBBO allows the use of surrogate probabilistic models tailored for combinatorial spaces with discrete variables. Any Bayesian model in which posterior inference is carried out through Markov chain Monte Carlo can be selected as the surrogate model in SBBO. We demonstrate empirically the effectiveness of SBBO using various choices of surrogate models in applications involving combinatorial optimization.
keywords:
Acquisition functions, Bayesian decision theory, Black-box functions, Combinatorial optimization, Markov chain Monte Carlo.
@@ 1 Introduction
The problem of optimizing a potentially nonlinear function $f(\boldsymbol{x})$ over a certain set $\mathcal{S}$ has been extensively studied, see e.g. [20], [16]. In many applications, the objective function is not explicitly known. This can make optimization challenging, as it is difficult to determine the optimal solution without accurate knowledge of the function. Moreover, $f(\boldsymbol{x})$ may be expensive to evaluate and a budget constraint often restricts evaluations of the objective function to a sparse selection of points, further complicating the optimization task. As an example, consider the process of drug discovery [10]. During its early stages, the goal is to identify candidate molecules with a desired property, such as maximum binding affinity to a given biological target. However, measuring this binding affinity requires synthesizing and testing molecules in vitro, limiting the number of leads that can be evaluated. Furthermore, expressing affinity as a mathematical function of the molecule structure is extremely difficult.
@@
Bayesian Optimization (BO, [21]) stands as an efficient approach for identifying the extrema of a black-box objective function. Its effectiveness originates from its capacity to integrate prior knowledge about the function with the evaluations previously conducted, enabling informed decisions on where to carry out additional evaluations. Bayesian Optimization involves the construction of a Bayesian surrogate model for $f(\boldsymbol{x})$. Subsequently, the next evaluation point is selected through the maximization of an acquisition function [9].
@@
Much of the BO literature focuses on continuous search spaces, where Gaussian Processes (GPs, [13]) are predominantly chosen as the Bayesian surrogate model. GPs are well-suited to optimization tasks in continuous search spaces as they rely on the smoothness defined by a kernel to model uncertainty. Additionally, GPs provide an analytical posterior predictive distribution, allowing for efficient computation of acquisition functions like expected improvement or probability of improvement. However, GPs are less effective in more complex scenarios, such as those involving combinatorial (categorical) or mixed categorical-continuous search spaces, where the variables include both discrete and continuous components. This paper specifically addresses the challenge of adapting BO to perform effectively in such complex search spaces.
Related work
In the last few years there has been some research in how to perform Bayesian optimization over combinatorial domains. As highlighted by [27], this task comes with two major challenges: constructing surrogate models that effectively capture the complexity of the combinatorial space, and efficiently searching within this space to identify the next candidate for evaluation using the acquisition function.
@@
Most previous research has been devoted to the first task. Within this, an important line of investigation consists of modifying GP kernel functions to accommodate categorical covariate spaces. For instance [11] treat categorical variables as continuous ones via one-hot encoding; then, within the kernel function, continuous covariates are suitably transformed to one-hot encodings. [25] construct a combinatorial graph over the search space; subsequently, graph Fourier transforms are used to derive a diffusion kernel on the graph. Another interesting line of research performs BO over a continuous latent representation of the combinatorial spaces learned through a Variational Autoencoder [12]. Finally, a third line of research gets rid of GPs, and proposes using inherently discrete surrogate models. For instance, [15] use random forests to model the objective function given the covariates. [2] model the response as a linear function of the covariates including all possible pairwise interactions. A heavy-tailed horseshoe prior is used to induce sparsity on the coefficients.
In this paper, we address the second key challenge in BO: efficiently searching within a combinatorial domain for the next candidate to evaluate after selecting an appropriate surrogate model. We introduce a novel simulation-based approach for optimizing acquisition functions in combinatorial domains, which we refer to as Simulation-Based Bayesian Optimization (SBBO). The key advantage of SBBO is its ability to optimize the acquisition function using only samples from the posterior predictive distribution, without requiring a closed-form expression. This makes BO applicable even when posterior inference is carried out through methods like Markov Chain Monte Carlo (MCMC). By removing the need for closed-form solutions, SBBO allows for the use of a broader range of surrogate probabilistic models that may be better suited for BO tasks in combinatorial spaces. This flexibility makes SBBO particularly effective in these contexts and significantly expands the applicability of BO.
@@
The paper is structured as follows. Section 2 provides an overview of Bayesian Optimization. Then, Section 3 presents SBBO and summarises the surrogate probabilistic models that we use for the experiments presented in Section 4. Finally, Section 5 gives some concluding remarks and lines of future research.
@@ 2 Bayesian Optimization
In the BO setting, the objective function $f(\boldsymbol{x})$ is assumed to be a black box, so no assumptions of, e.g., convexity or linearity are made. Additionally, it is assumed that $f$ is expensive to evaluate in that the number of evaluations of $f$ that can be feasibly performed is small, perhaps due to computing time or financial cost. It is further assumed that querying the function at $x$ provides a noisy observation $y$ of $f(x)$.
BO uses a surrogate probabilistic model of the objective function, usually obtained through Bayesian methods, to inform the next evaluation decision. It works iteratively by: 1) fitting the surrogate model, and 2) using an acquisition function to obtain a new point $\boldsymbol{x}$ at which we evaluate or observe $f(\boldsymbol{x})$. The surrogate model is typically fitted using all available data $\mathcal{D}_{1:n}=\{(\boldsymbol{x}_{i},y_{i}),i=1,\ldots,n\}$. After selecting the new evaluation location $\boldsymbol{x}_{n+1}$, the corresponding $y_{n+1}$ is obtained, and the surrogate model is updated to include the additional data point $(\boldsymbol{x}_{n+1},y_{n+1})$. Let us briefly review these two key components of Bayesian Optimization.
For the surrogate model, the following observational equation is assumed
@@
where $f(\boldsymbol{x})$ could be parametric or non-parametric. Given a new location $\boldsymbol{x}$ and appropriate choice of prior, the posterior predictive distribution on $f(\boldsymbol{x})$ is taken as the surrogate model. If a parametric model is chosen for $f$, then $f(\boldsymbol{x})$ will depend on parameters $\boldsymbol{\beta}$ and is denoted as $f_{\boldsymbol{\beta}}(\boldsymbol{x})$. A prior $\pi(\boldsymbol{\beta})$ for the parameters $\boldsymbol{\beta}$ needs to be specified. Coupled with a likelihood $\pi(y|\boldsymbol{x},\boldsymbol{\beta},\sigma^{2})$, after observing data $\mathcal{D}_{1:n}$, we obtain the posterior $\pi(\boldsymbol{\beta}|\mathcal{D}_{1:n})$. Then, the surrogate model is
$\pi(f(\boldsymbol{x})|\boldsymbol{x},\mathcal{D}_{1:n})\equiv\pi(f_{\boldsymbol{\beta}}(\boldsymbol{x})|\boldsymbol{x},\mathcal{D}_{1:n})=\mathbb{E}_{\boldsymbol{\beta}|\mathcal{D}_{1:n}}\left[f_{\boldsymbol{\beta}}(\boldsymbol{x})\right]$ (2.3)
Note that, in most cases, the expectation in (2.3) does not have an analytical form. Typically, the only option is to generate samples from the surrogate model by first sampling $\beta\sim\pi(\boldsymbol{\beta}|\mathcal{D}_{1:n})$, usually via MCMC, and then computing $f_{\boldsymbol{\beta}}(\boldsymbol{x})$.
@@
Regarding the acquisition function, there are several possible choices. We restrict our attention to acquisition functions that can be written as expectations of a utility function with respect to the posterior $\pi(f(\boldsymbol{x})|\boldsymbol{x},\mathcal{D}_{1:n})$. Some of the most commonly used acquisition functions belong to this class, such as the expected improvement and the probability of improvement. Let $u(f(\boldsymbol{x}),\boldsymbol{x})$ denote the utility achieved by evaluating the function at a new point $\boldsymbol{x}$ and obtaining $f(\boldsymbol{x})$. Then, the next point $\boldsymbol{x}_{n+1}$ is chosen as the one that maximizes the acquisition function, i.e. the expected utility:
$\Psi(\boldsymbol{x})\equiv\mathbb{E}_{f|\boldsymbol{x},\mathcal{D}_{1:n}}\left[u(f(\boldsymbol{x}),\boldsymbol{x})\right].$ (2.4)
When using a GP prior for $f$, the posterior distribution is Gaussian, and for common utility functions, there is a closed-form expression for the expectation in Equation (2.4) that facilitates its optimization. The BO framework is flexible, but its computational efficiency and analytical results rely on this Gaussian process prior. However, in specific scenarios, such as when the search space includes categorical or mixed categorical-continuous variables, a GP prior may not be the most suitable choice. In these cases, alternative Bayesian models better suited to the problem may be available. However, these alternatives often lack an analytically-defined posterior distribution, complicating the selection of the next point $\boldsymbol{x}_{n+1}$ based on Equation (2.4). Nonetheless, for most Bayesian models, we can sample from the posterior predictive distribution via MCMC methods. To address this challenge when employing a non-GP prior, we introduce a Bayesian optimization methodology that exclusively requires sampling access to the posterior. This approach enables the use of Bayesian models that are better suited to structured covariate spaces.
It is important to emphasize that choosing the next sampling location in (2.4) implicitly entails adopting a “one-step” assumption: the selection is made under the presumption that no future evaluations will take place. When there is budget for more than one function evaluation, expected improvement and probability of improvement optimization unequivocally adopt a myopic approach. While several alternative strategies have been proposed for multi-step lookahead Bayesian Optimization [29, 18], we focus on the one-step look-ahead case.
@@ 3 Simulation Based Bayesian Optimization
We adapt ideas from the simulation-based optimal Bayesian design literature [23] into the field of Bayesian Optimization, yielding our proposed method of Simulation-Based Bayesian Optimization (SBBO).
@@ 3.1 Main Idea
We aim to solve the optimization problem 2.1 with a limited number of function evaluations. Suppose we have access to data $\mathcal{D}_{1:n}=\{(\boldsymbol{x}_{i},y_{i}),i=1,\ldots,n\}$, possibly obtained from previous experiments, where $\boldsymbol{x}_{i}$ is a vector of covariates for observation $i$ and $y_{i}$ is derived from the observation model (2.2). Given a budget of at most $M$ additional function evaluations, our task is to determine the next evaluation points $\boldsymbol{x}_{n+1},\boldsymbol{x}_{n+2},\ldots,\boldsymbol{x}_{n+M}$. As introduced in Section 2, these points are usually selected in a greedy manner. Let $u(f(\boldsymbol{x}),\boldsymbol{x})$ be a non-negative and bounded utility function that quantifies the benefits or gains from evaluating the objective function at $\boldsymbol{x}$ and observing the corresponding $f(\boldsymbol{x})$. Following Bayesian decision-theoretic principles, the next evaluation should be performed at the location $\boldsymbol{x^{*}}$ that maximizes the expected utility (2.4).
@@
To perform Bayesian optimization, we need to select a suitable Bayesian surrogate model and optimize (2.4). As mentioned earlier, except in a few special cases, the posterior $\pi(f(\boldsymbol{x})|\boldsymbol{x},\mathcal{D}_{1:n})$ is generally only accessible through sampling, typically using MCMC methods. The specifics of the sampling process depend on the surrogate model employed (see Section 3.3 for examples). Assuming we can generate samples from $\pi(f(\boldsymbol{x})|\boldsymbol{x},\mathcal{D}_{1:n})$, we can borrow ideas from Bayesian decision analysis and reframe the optimization of (2.4) as a simulation problem. Following [4], we define an augmented distribution $g$ that treats $\boldsymbol{x}$ as a random variable:
$g(\boldsymbol{x},f(\boldsymbol{x})|\mathcal{D}_{1:n})\propto u(f(\boldsymbol{x}),\boldsymbol{x})\pi(f(\boldsymbol{x})|\boldsymbol{x},\mathcal{D}_{1:n}).$
@@
It is straightforward to see that the marginal of $g$ in $\boldsymbol{x}$ is proportional to the expected utility (2.4). Thus, the mode of this marginal distribution coincides with the optimal next evaluation point. This suggests a simulation-based approach to find $\boldsymbol{x}_{n+1}$: we simulate $(\boldsymbol{x},f(\boldsymbol{x}))\sim g(\boldsymbol{x},f(\boldsymbol{x})|\mathcal{D}_{1:n})$ using standard MCMC techniques and compute the mode of $g(\boldsymbol{x}|\mathcal{D}_{1:n})$ using the generated samples $\boldsymbol{x}$. However, this strategy is limited to low-dimensional $\boldsymbol{x}$. For high-dimensional covariate spaces, estimating the mode of $g(\boldsymbol{x}|\mathcal{D}_{1:n})$ from the simulation output becomes impracticable. [24] propose an alternative augmented distribution $g_{H}$ defined as
@@
where $f(\boldsymbol{x})_{1},\ldots,f(\boldsymbol{x})_{H}$ are $H$ copies of the random variable $f(\boldsymbol{x})$. Now, as the marginal of $g_{H}$ in $\boldsymbol{x}$ is
$g_{H}(\boldsymbol{x}|\mathcal{D}_{1:n})\propto\Psi(\boldsymbol{x})^{H},$
the mode of this distribution also coincides with the optimal point at which we should make our next function evaluation. In particular, if $H$ is sufficiently large, the marginal distribution $g_{H}(\boldsymbol{x})$ tightly concentrates on the next $\boldsymbol{x}_{n+1}$, thus facilitating mode identification.
[24] propose an inhomogeneous MCMC sampling scheme to simulate from $g_{H}$. Their approach begins by generating samples $(\boldsymbol{x},f(\boldsymbol{x}))\sim g(\boldsymbol{x},f(\boldsymbol{x})|\mathcal{D}_{1:n})$ using the standard Metropolis-Hastings algorithm. In subsequent iterations, the target distribution is adjusted to $g_{H}$, with $H$ increasing incrementally. This inhomogeneous MCMC is designed so that the stationary distribution for a fixed $H$ is $g_{H}$. We build upon this idea by adapting it to find the next evaluation point in the context of Bayesian optimization. Our proposed approach is detailed in Algorithm 1, where we utilize a Metropolis-Hastings sampling scheme. For clarity, we choose the state vector of the Markov chain to be $(\mathbf{x},v)$ rather than $(\mathbf{x},f(\mathbf{x})_{1},\dots,f(\mathbf{x})_{H})$, where $v=\frac{1}{H}\sum_{h=1}^{H}\log u(f(\mathbf{x})_{h},\mathbf{x})$.
@@
Obtain initial dataset $\mathcal{D}_{1:n}=\{(\boldsymbol{x}_{i},y_{i}),i=1,\ldots,n\}$. Determine utility function $u_{\mathcal{D}}(\boldsymbol{x},y)$.
@@
    Choose initial $\boldsymbol{x}_{n+1}^{(0)}$. For $h=1,\ldots,H$, simulate $f(\boldsymbol{x})^{(0)}_{h}\sim\pi(f(\boldsymbol{x})|\boldsymbol{x}_{n+1}^{(0)},\mathbf{D}_{1:n})$, evaluate utility $u^{(0)}_{h}=u(f(\boldsymbol{x}_{n+1}^{(0)})_{h},\boldsymbol{x}_{n+1}^{(0)})$. Define $v^{(0)}=\frac{1}{H}\sum_{h=1}^{H}\log u^{(0)}_{h}$.
@@
    For $h=1,\ldots,H$, simulate $\tilde{f}(\boldsymbol{x})_{h}\sim\pi(f(\boldsymbol{x})|\tilde{\boldsymbol{x}}_{n+1},\mathcal{D}_{1:n})$ and evaluate utility $\tilde{u}_{h}=u(\tilde{f}(\boldsymbol{x})_{h},\tilde{\boldsymbol{x}}_{n+1})$. Define $\tilde{v}=\frac{1}{H}\sum_{h=1}^{H}\log\tilde{u}_{h}$.
@@
    $\alpha$ $=\min\left\{1,\exp\left(H\tilde{v}-Hv^{(0)}\right)\cdot\frac{q(\boldsymbol{x}^{(0)}_{n+1}|\tilde{\boldsymbol{x}}_{n+1})}{q(\tilde{\boldsymbol{x}}_{n+1}|\boldsymbol{x}^{(0)}_{n+1})}\right\}$